A monopoly firm faces a demand curve given by the following equation: P = $500 − 10Q, where Q

equals quantity sold per day. Its marginal cost curve is MC = $100 per day. Assume that the firm faces

no fixed cost. You may wish to arrive at the answers mathematically, or by using a graph (the graph is

not required to be presented), either way, please provide a brief description of how you arrived at your

results.



1
Expert's answer
2021-12-02T10:38:42-0500

Solution:

A monopoly maximizes profit where MR = MC

Derive MR:

TR = P "\times" Q = (500 – 10Q) "\times" Q = 500Q – Q2

MR = "\frac{\partial TR} {\partial Q}" = 500 – 2Q

 

Set MR = MC:

500 – 2Q = 100

500 – 100 = 2Q

400 = 2Q

Q = 200

Profit maximizing quantity = 200

Substitute in the demand function to derive price:

P = 500 – Q = 500 – 200 = 300

Profit maximizing price = 300